OpenStudy (yinshy):

Which is the best approximation for the circumference of a circular fish pond that is 24 ft in diameter? A. 38 ft B. 72 ft C. 75 ft D. 150 ft

OpenStudy (yinshy):

@MysticFury

rishavraj (rishavraj):

r = d/2 = 12 circumference = 2*pi*r

OpenStudy (mysticfury):

looks like @rishavraj beat me to it lol

OpenStudy (richhomiehelp):

sup @Yinshy

rebeccaxhawaii (rebeccaxhawaii):



OpenStudy (mysticfury):

or @Yinshy you could just times the diameter by 3.14 since the radius is timesed by 2 anyway

OpenStudy (mysticfury):

so what do you think is the answer @Yinshy ?

OpenStudy (yinshy):

C